Water flows over Victoria Falls, which is 128 m high, at an average rate of 1.39 × 106 kg/s. The acceleration of gravity is 9.81 m/s 2 . If half the potential energy of this water were converted into electric energy, how much power would be produced by these falls? Answer in units of MW.
Answer #1

Potential energy of water is PE=mgh

Power is the rate at which the work is done , P=rac{mgh}{t}=rac{m}{t}gh

Given that, Height from where water falls is 11 128 m

Rate at water falls in time t is 1.39 106 kg/s

Half of the potential energy is converted into electric energy,

Power produced by water falls is  rac{P}{2}=rac{m}{t}rac{gh}{2}=1.39*10^6*rac{9.81*128}{2}=872.7*10^6,W=872.6,MW